Find the solutions of the equation
x4 − 4x3 − x2 + 4x = 0.
(Enter your answers as a comma-separated list.)

Answer by stanbon(75887)

Find the solutions of the equation
x4 − 4x3 − x2 + 4x = 0.
------------------
x(x^3 - 4x^2 - x + 4) = 0
-------
x[x^2(x-4)-(x-4)] = 0
-------
x(x-4)(x^2-1) = 0
x(x-4)(x-1)(x+1) = 0
-----------------------
x = 0 or x = 4 or x = 1 or x = -1
-----------------------------------
Cheers,
Stan H.
-----------------